OpenStudy (milo123):

MEDAL I HAVE 5 MINUTES TO TURN THIS IN (IS IT B?) The leg of a right triangle is 3 units and the hypotenuse is 4 units. What is the length, in units, of the other leg of the triangle? Square root of 5 Square root of 7 5 7

OpenStudy (milo123):

is it B?

OpenStudy (wolf1728):

leg^2 = 4^2 - 3^2 leg^2 = 16 -9 leg = sq root (7)

OpenStudy (wolf1728):

milo123 Did you get it right?